ZAKOYA7WAITERS:

On your own: A blue whale has a length of 28,578 millimeters. What is the length of a blue whale in kilometers? Show your work below by writing the conversion factors you use to solve this question. 28,578 mm • ____________ • ____________ = 0.028578

dude:

You must know conversions 1mm=0.001m 1m=0.001km

kittybasil:

what dude said, & also your format should be:\[\frac{28,578\text{ nm}}{1}\cdot\frac{\text{?? m}}{\text{?? mm}}\cdot\frac{\text{?? km}}{\text{?? m}}\] Note that you can cross cancel some of the units...

kittybasil:

*sorry; there is no "nm" ... that is "mm" but it's a typo

dude:

Its nm :) ;o